Operations Report Jan 18th

Crew 203 Operations Report 18-JAN-2019

SOL: 6

Name of person filing report: Freddy Castaneda

Non-nominal systems: Nothing new to report.

Notes on non-nominal systems: None

Generator (hours run): 6hr 40min; Turned on last night (17Jan2019) at 18:15; The generator shuted down at 20:40 and did not start again. All electrical support over the night was provided by the batteries of the SOC. The generator was turned on this morning (18Jan2019) at 9:25; It was turned of again at 13:40 after SOC batteries were full charged and turned on tonight (18Jan2019) at 16:29

Solar SOC – Turned on (17Jan2019) 74%; It wasn’t turned off (18Jan2019) until it was full charged at 13:40, at 9:30 it was on 38% of charge, at 10:30 it was on 48% and at 13:40 it was ful of charge 100%; Turned on (19Jan2019) 100%

Diesel Reading – 1/2 tank.

Propane Reading – Just over 25%

Ethanol Free Gasoline – Not in use

Water (auxiliary tank) – Not in use

Water (static tank) – Under 70%; About 350 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 70%; 211 gallons

Water (loft) – At level marker 10

Static to Loft Pump used – Yes; At 10:20 to refill tank

Water Meter: 014012054

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Not used

Hours: 68.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Opportunity rover used: Not used

Hours: 45.9

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Not used

Hours: 70.0

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: None

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: None

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? Not used.

CrewCar used and why, where? Off campus

General notes and comments: The generator was shuted down all night, it was fixed by Shannon and Atila on the morning.

Summary of internet: Nothing to report

Summary of suits and radios: Suit #3 still pending for review, Suit #8 charged normally and is functional.

Summary of Hab operations: Nothing to report.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: Nothing to report

Freddy